OpenStudy (loser66):

Prove: If L is a line in \(R^2\) and if T: \(R^2 \rightarrow R^2\) is a linear transformation, then T(L) is also a line in \(R^2\) Please, help

OpenStudy (loser66):

@perl

OpenStudy (perl):

we have to use the definition of linear transformation

OpenStudy (perl):

do you have the book of this online

OpenStudy (loser66):

I have definition of linear transformation from the book I took the course last semester. hihihi. Let me write it out

OpenStudy (perl):

it is blurry, can you do the pic again

OpenStudy (perl):

a line in R^2 has the form ax +by = c

OpenStudy (loser66):

I am sorry, my scanner is stupid, let me type it down

OpenStudy (loser66):

A linear transformation (or operator) A on a vector space V is a correspondence that assigns to every vector x in V a vector Ax in V, in such a way that \[A(\alpha x+\beta y)= \alpha Ax +\beta A y\]

OpenStudy (loser66):

We can do something like: Let \(\vec x=\left [\begin {matrix}x_1\\x_2\end{matrix}\right]\) and \(\vec y=\left [\begin {matrix}y_1\\y_2\end{matrix}\right]\) in \(R^2\)

OpenStudy (loser66):

\(\alpha\), \(\beta\) are scalar Since T is linear transformation , we have \(T(\alpha\vec x)=T\left [\begin {matrix}\alpha x_1\\\alpha x_2\end{matrix}\right]\) and \(T(\beta \vec y)=T\left [\begin {matrix}\beta y_1\\\beta y_2\end{matrix}\right]\)

OpenStudy (loser66):

and then?? hihihi

OpenStudy (loser66):

oh, and then the distance between x and y = distance between Tx and Ty. Is it valid??

OpenStudy (perl):

well you want a line

OpenStudy (perl):

how we do we describe a line in vector notation r(t) = ro + t*V

OpenStudy (loser66):

no idea!! :(

OpenStudy (xapproachesinfinity):

@Loser66 are you a linear algebra fan hehe! when i took this course, i was so lost haha

OpenStudy (xapproachesinfinity):

well i actually passed it, i would say it was B minus in american system of grading

OpenStudy (xapproachesinfinity):

oh i see you are still better anyway, so why you say shame on you haha

OpenStudy (xapproachesinfinity):

bad scanner bro haha

OpenStudy (kirbykirby):

This is slightly distant in my memory.. but if your line \(L\) is written as \(\vec{x}=\alpha\vec{m}+\vec{b}, \alpha\in \mathbb{R}\), then \(T(L)=T(\alpha\vec{m}+\vec{b})=\alpha T(\vec{m})+T(\vec{b})\) since \(T\) is a linear mapping. And since \(T: \mathbb{R}^2\rightarrow \mathbb{R}^2\) , then \(T(\vec{m})\in \mathbb{R}^2 \) and \(T(\vec{b})\in \mathbb{R}^2\) , and so you can think of the new line being like \(\alpha \vec{v}+\vec{w}\) where \(\vec{v}=T(\vec{m})\) and \(\vec{w} =T(\vec{b})\) I'm not 100% sure though.

OpenStudy (loser66):

I got all but the very first expression \[\vec{x}=\alpha\vec{m}+\vec{b}, \alpha\in \mathbb{R}\] How can we write the line under vector form like this?

OpenStudy (perl):

that is the vector form

OpenStudy (loser66):

Got it, hihihi

OpenStudy (loser66):

Thank you for the help, friends. I appreciate. :)

OpenStudy (perl):

yes it looks like the proof works for any line in R^n

OpenStudy (perl):